OpenStudy (anonymous):

Find the midpoint of the segment with endpoints A(-7,6) and B (-1,1)

OpenStudy (anonymous):

use the midpoint formula \[\left( \frac{x _{1 + X _{2} } }{ 2 } \right) , \left( \frac{y _{1 +Y _{2} } }{ 2 } \right)\]

OpenStudy (anonymous):

\[x _{1} \] is -7 , \[x _{2} \] is -1 . \[y _{1} \] is 6 , \[y _{2} \] is 1 .